OpenStudy (anonymous):

solve the inequality -1/2y+6<0

OpenStudy (anonymous):

you would subtract by 6 on both sides then multiply by 2 and switch the sign i believe

OpenStudy (tommynaut):

I'll give you a similar example with working out. \[-5 -\frac{ 1 }{ 4 } y > 3\] We solve this just like how we'd solve a normal equation, at first. We want to isolate the y term, so what else is there apart from the y term on the left hand side? The -5! To get rid of negative 5, we add 5, because -5 +5 = 0. So, we +5 to both sides. \[-5 + 5 - \frac{ 1 }{ 4 }y > 3 + 5\] Which simplifies to \[-\frac{ 1 }{ 4 }y > 8\] Now, we want y COMPLETELY BY ITSELF. So, the last step is to get rid of the number next to it. We have y being multiplied by negative 1 quarter, so to get rid of it, we divide both sides by negative 1 quarter. However, this is the weird part. If you ever multiply or divide both sides of an inequality (like, > or <) then you have to flip the inequality sign. So we'll get: \[-\frac{ 1 }{ 4 }y \div -\frac{ 1 }{ 4 } < 8 \div -\frac{ 1 }{ 4 }\] \[y < -32\] See if you can apply the same method to your question!